The orbital speed of electron orbiting around the nucleus in a circular orbit of radius 50Pm is 2.2x 106 m/s, the magnetic dipole moment of an electron is
1
Expert's answer
2020-06-24T11:22:05-0400

As per the given question,

"e=1.6\times 10^{-19}Coulomb"

The radius of the circular orbit (r)=50pm

Speed of electron "(v)=2.2 \times 10^6 m/sec"

magnetic dipole moment of an electron =?

Magnetic dipole moment "(m)=iA=\frac{e}{T}\pi r^2"


"=\frac{e}{\frac{2\pi r}{v}}\pi r^2 =\frac{erv}{2}"


"=\frac{1.6\times 10^{-19}\times50\times 10^{-12}\times 2.2 \times 10^6}{2} \times \frac{Coulomb\times m\times m}{sec}"


"=8.8\times 10^{-24}\frac{Coulomb}{sec}\times m^2"


"=8.8\times 10^{-24}Am^2" (because we know that "Ampere=\frac{Coulomb}{sec}" )